Two forces, F⃗ 1 and F⃗ 2, act at a point. F⃗ 1 has a magnitude of 8.00 N and is directed at an angle of 65.0 ∘ above the negative x axis in the second quadrant. F⃗ 2 has a magnitude of 7.00 N and is directed at an angle of 53.9 ∘ below the negative x axis in the third quadrant.

Part A

What is the x component of the resultant force?

Express your answer in newtons.

N

Part B

What is the y component of the resultant force?

Express your answer in newtons.

N

Part C

What is the magnitude of the resultant force?

Express your answer in newtons.
Answer #1

here,

F1 = 8 * (- cos(65) i + sin(65) j ) N

F1 = - 3.38 i N + 7.25 j N

F2 = 7 * ( - cos(53.90) i + sin(53.9) j ) N

F2 = - 4.12 N i - 5.66 j N

A)

the x component of resultant force , Fx = F1x + F2x

Fx = - 7.5 N

B)

the y component of resultant force , Fy = F1y + F2y

Fy = + 1.59 N

C)

F = sqrt(Fx^2 + Fy^2)

F = 7.67 N